OpenStudy (anonymous):

The fifth and the seventh terms of a geometric sequence are 3 and 9. What are all POSSIBLE values of the sixth term of the sequence?

OpenStudy (anonymous):

a_5 = ar^4 = 3 a_6 = ar^5 = 3r a_7 = ar^6 = 3r^2 = 9 -> r = plus minus sqrt(3) and a_6 = plus minus 3 sqrt 3